OpenStudy (anonymous):

Find the length of the hypotenuse of a 45°- 45°-90° triangle with a leg of 8.

OpenStudy (raden):

the proportion sides of a 45°- 45°-90° triangle is 1 : 1 : sqrt(2)

OpenStudy (anonymous):

so is that the answer or do you have to proportion it according to the equation?

OpenStudy (raden):

no, that's just a hint... if given a leg of right isosceles triangle is x, so the length of hypotenusa is xsqrt(2)

OpenStudy (anonymous):

so this one would be [8sqrt{2}] ?

OpenStudy (raden):

yes, correct

OpenStudy (anonymous):

ok now i have a nother one that i need help with A square piece of paper 140 mm on a side is folded in half along a diagonal. The result is a 45° – 45° – 90° triangle. What is the length of the hypotenuse?

OpenStudy (raden):

do same idea like number 1

OpenStudy (anonymous):

so would the sides on this triangle stay the same.. 140mm for the two legs?

OpenStudy (raden):

yes, it would be a right isosceles triangle...

OpenStudy (anonymous):

140sqrt(2) ?

OpenStudy (raden):

yes

OpenStudy (anonymous):

i need to simplify sqrt(200)

OpenStudy (raden):

sqrt(200) = sqrt(100*2) = sqrt(100) * sqrt(2) = 10sqrt(2)

OpenStudy (anonymous):

so simplifying sqrt(20) would be basicly the same as doin that?

OpenStudy (raden):

yes

OpenStudy (anonymous):

ok got it. do you know triangle inequalities too?

OpenStudy (raden):

i will try it

OpenStudy (anonymous):

in triangle PQR line PQ= 7.2 and line QR= 5.2. Which measure cannot be PR? 7,9,11,13

OpenStudy (raden):

7.2 + 5.2 > 13 12.4 > 13 (a mistake statement) so, 13 the answer